OpenStudy (anonymous):

Is this linear or nonlinear? \[y(t) = \frac{2x(t)}{2x(t) -1}\]

OpenStudy (anonymous):

linear

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Because when you simplify it, it is in slope-intercept form

OpenStudy (anonymous):

and if the power of the variable is then it is also linear

OpenStudy (anonymous):

it's rational though

OpenStudy (anonymous):

how does it simplify to slope-intercept form anyway?

OpenStudy (anonymous):

did you simplify it correctly so that it is in slope intercept form?

OpenStudy (anonymous):

i see nothing that can simplify it further. I can't even turn it into partial fractions

OpenStudy (amistre64):

there are a few definitions of linear

OpenStudy (amistre64):

a function is linear if: f(a+b) = f(a) + f(b) and f(ka) = k f(a) for some constant k

OpenStudy (anonymous):

would it make a difference if I actually said this is a signal function and not just ordinary algebraic function? I don't think it does, but does it?

OpenStudy (amistre64):

it might help :) can you define what a signal function is?

OpenStudy (anonymous):

i don't know its exact definition but I think it's something that is a function of time

OpenStudy (phi):

you could use amistre's conditions to test your function: for example, evaluate it at x=2 then evaluate it at x= 3*2 if it is linear f( k x) = k f(x) which means (with k=3) f(6) = 3 * f(2) btw, it won't be

OpenStudy (amistre64):

i cant find a suitable reference thats readable from the google to help clarify it either

OpenStudy (amistre64):

from the looks of setup, it appears to be a rational function what does not conform into a linear or affine nature. 2x-3y = 0 is a linear function 3x+4y-7z = 0 is a linear function affine functions tend to be constructed as: {linear} + C, for some constant C. And are viewd by many to be "linear"

OpenStudy (anonymous):

but even in algebra, i don't think rational functions can be linear right?

OpenStudy (phi):

generally true